Inverse of x²+5x+9
Found 2 solutions by Boreal, natolino_2017:
Answer by Boreal(15235)

y=x^2+5x+9
x=y^2+5y+9
y^2+5y=x-9
y^2+5y+6.25=x-9+6.25, completing the square
(y+2.5)^2=x-2.75
y=-2.5+/- sqrt(x-2.75)


Answer by natolino_2017(77)

let f(x) = x^2 + 5x + 9.
domain of f: all the real numbers.
f(-3) = 3 and f(-2) = 3, so the function is not injective.
So there's not an inverse for the function for the domain.
@natolino